For the past year, Allargando Company recorded sales of $496,800 and average operating assets of $276,000. What is the margin that Allargando Company needed to earn in order to achieve an ROI of 12.6%?

7.94%

1.8%

7%

12.6%
Answer #1

ROI=net operating income/average operating assets

net operating income=276,000*12.6%=34776

Margin=net operating income/Sales

=34776/496,800

=7%